# GROOD: Gradient-Aware Out-of-Distribution Detection

### v2

@@ 6.1 Ablation Studies
To understand the contribution of each component in GROOD, we conduct a series of ablation studies on CIFAR-10 (table 4). Each study isolates a key design choice and quantifies its impact on both near-OOD and far-OOD detection performance.
@@ 6.2 Robustness to Checkpoint Choice
The AUROC metric, while widely used for evaluating OOD detection, can exhibit instability during training, particularly in the later stages. This instability means that small fluctuations in the model’s weights can lead to significant variations in AUROC scores, making the selection of an optimal checkpoint challenging. The AUROC curve can vary sharply, even when the test error is relatively stable, indicating a sensitivity to minor weight perturbations. In contrast, GROOD’s design contributes to more stable OOD detection performance. The key intuition behind GROOD’s robustness lies in its focus on the sensitivity of weights relative to the OOD prototype. Throughout training, while the representation space and the OOD prototype’s absolute location change as the network’s weights are updated, their inherent relationship—the sensitivity—remains stable, leading to consistent OOD detection , as further evidenced by the reduced standard deviations shown in table 5 and fig. 3.
Further ablation experiments and detailed analyses are provided in the Appendix.
@@ 7 Discussion and Conclusion
In this work, we presented GRadient-aware Out-Of-Distribution detection (GROOD), a novel approach to address the critical challenge of detecting OOD samples in DNN-based image classifiers. GROOD leverages the complementary strengths of gradient information and distance metrics, demonstrating a significant advancement in identifying OOD instances, which is crucial for deploying DNNs reliably and safely in real-world applications (Litjens et al., 2017; Bojarski et al., 2016). The extensive experimental evaluation across various benchmarks illustrates the efficacy of GROOD in detecting both near and far OOD samples, exhibiting robust performance independent of specific DNN architectures or training data distributions. Furthermore, GROOD’s ability to operate effectively without extensive hyper-parameter tuning underscores its practicality and adaptability in diverse settings.
However, it is important to consider the appropriate use cases and potential limitations of GROOD. As GROOD relies on the concept of class prototypes and is inspired by the Neural Collapse phenomenon (Papyan et al., 2020), it is inherently well-suited for applications where the in-distribution data exhibits a clear cluster structure in the feature space. In such scenarios, GROOD’s ability to effectively distinguish between these clusters and identify deviations becomes particularly advantageous. Conversely, GROOD may face challenges in applications where the in-distribution data is inherently noisy, or where the underlying model fails to clearly separates the in-distribution classes.
Beyond application-specific considerations, GROOD also presents certain practical limitations. While our approach demonstrates improved inference speed compared to KNN-based methods, as detailed in § A.4, there are computational costs associated with storing and processing gradient information, which can become significant for very large datasets or high-resolution images. Furthermore, the construction of effective OOD prototypes requires careful consideration. As discussed in § 5.2, while GROOD is relatively robust to the specific choice of auxiliary OOD samples, the diversity and representativeness of these samples can influence performance as shown in § A.2.
Several subtle but important design choices were made during the development of GROOD. For example, the proximity-based filtering step in OOD prototype construction, as described in § 5.2, proved crucial for enhancing the discriminative power of the prototype. We also observed that using a mixup strategy targeting the second-highest predicted class for synthetic OOD data generation led to better results than interpolating towards a random class, likely due to the resulting samples being closer to the decision boundary.
In conclusion, GROOD offers a promising approach for OOD detection, particularly in applications with well-structured in-distribution data. Future research could explore methods to extend GROOD to handle more complex data distributions or to further optimize its computational efficiency.
@@ A.1 GROOD Algorithm
For a comprehensive overview, the complete GROOD algorithm is outlined in algorithm 1 and algorithm 2.
@@
3:Compute class prototypes $\mathbf{P}_{\text{early}}$ and $\mathbf{P}_{\text{pen}}$triangleright eq. 7
4:Compute OOD prototype $\mathbf{p}^{\text{pen}}_{\text{ood}}$ using synthetic data generation§ 5.2 $\triangleright$ eq. 8
@@
6:     compute $\nabla H(h)$triangleright eq. 5 using $\mathbf{p}^{\text{pen}}_{y=1,\cdots,C}$, $\mathbf{p}^{\text{pen}}_{\text{ood}}$
@@ A.2 Choice of OOD data for OOD prototype computation
To simulate OOD data representations, we utilize a manifold mixup technique on the early layer, similar to the targeted mixup approach described in § 5. However, our method differs in the interpolation target. Instead of interpolating towards the predicted class prototype, we interpolate towards the second-highest predicted class $c_{2}$, which is the closest incorrect class on the decision boundary.
@@
Table 6 illustrates our method’s robustness to different validation OOD datasets.
@@ A.3 Density Plots
To comprehensively evaluate the GROOD method’s ability to distinguish between in-distribution (ID) and out-of-distribution (OOD) data, we visualize the distribution of OOD scores across a range of datasets with varying characteristics. Figure 4 presents these visualizations for ID, Near-OOD, and Far-OOD samples on CIFAR-10, CIFAR-100 (datasets with relatively small, natural images), ImageNet-200 (a subset of ImageNet), and ImageNet-1k (a large-scale, complex dataset). Analyzing performance across this spectrum demonstrates GROOD’s robustness to differences in image complexity and dataset size. In each subplot, we use density plots to represent the distribution of OOD scores, allowing for a clear visual comparison of separation between ID and OOD distributions.
@@ A.4 Inference Speed
GROOD introduces a novel Out-of-Distribution (OOD) detection method involving two forward passes for the mixup part which is inexpensive to compute, a backward pass over the OOD prototype which can be computed using its closed form expression as in eq. 5 of the main paper and the ne arest neighbor search which is more computationally intensive. For the latter, Our approach employs the FAISS IndexIVF method for efficient distance computation, utilizing centroids and inverted lists instead of the complete dataset. This technique notably enhances inference speed compared to KNN, particularly in our CIFAR benchmarks. Specifically, on CIFAR-10 and CIFAR-100 datasets, GROOD recorded evaluation inference times over all OOD test sets of 130 seconds and 155 seconds, respectively. This is significantly faster than KNN, which took 434 seconds for CIFAR-10 and 641 seconds for CIFAR-100, demonstrating the efficiency of our approach.
@@ A.5 Ablate value of kth nearest neighbor
Figure 5 illustrates the impact of using the distance to the $k$-th nearest neighbor, as proposed by Sun et al. (2022). The plot demonstrates that employing the distance to the nearest point in gradient space leads to optimal results.
@@ A.6 Gradient computation in closed form
This section details the derivation of the closed-form expression for the gradient, as presented in eq. 5 of the main paper. The cross-entropy loss for a logit $L=(L_{j})_{j=1}^{C}+1$ and some (any) ID label $y$ is given by:
@@
which is equal to the softmax probability corresponding to the OOD class and does not depend on the specific ID class label $y$ anymore. Now for a feature vector $h$, the corresponding logit vector $L(h)$ is given by eq. 2. Since $L_{C+1}(h)=\|h-\mathbf{p}^{\text{pen}}_{\text{ood}}\|_{2}$ is the only logit depending on $\mathbf{p}^{\text{pen}}_{\text{ood}}$, the gradient of the above loss with respect to $\mathbf{p}^{\text{pen}}_{\text{ood}}$ is given by the chain rule:

### v3

@@ 6.1 Ablation Studies
To understand the contribution of each component in GROOD, we conduct a series of ablation studies on CIFAR-10 (table˜4). Each study isolates a key design choice and quantifies its impact on both near-OOD and far-OOD detection performance.
@@ 6.2 Robustness to Checkpoint Choice
The AUROC metric, while widely used for evaluating OOD detection, can exhibit instability during training, particularly in the later stages. This instability means that small fluctuations in the model’s weights can lead to significant variations in AUROC scores, making the selection of an optimal checkpoint challenging. The AUROC curve can vary sharply, even when the test error is relatively stable, indicating a sensitivity to minor weight perturbations. In contrast, GROOD’s design contributes to more stable OOD detection performance. The key intuition behind GROOD’s robustness lies in its focus on the sensitivity of weights relative to the OOD prototype. Throughout training, while the representation space and the OOD prototype’s absolute location change as the network’s weights are updated, their inherent relationship the sensitivity remains stable, leading to consistent OOD detection , as further evidenced by the reduced standard deviations shown in table˜5 and fig.˜3.
Further ablation experiments and detailed analysis are provided in the Appendix.
@@ 7 Discussion and Conclusion
In this work, we introduced GRadient-aware Out-Of-Distribution detection (GROOD), a novel approach that combines gradient information with distance metrics to improve detection of OOD samples in DNN-based image classifiers. Extensive experiments across benchmarks show that GROOD effectively detects both near and far OOD samples, performs robustly across architectures and datasets, and requires little hyper-parameter tuning, making it practical for deployment in real-world settings (Litjens et al., 2017; Bojarski et al., 2016).
GROOD is particularly well-suited to scenarios where in-distribution features form clear clusters, consistent with the Neural Collapse phenomenon (Papyan et al., 2020). In such cases, deviations from prototypes are easily detected. However, its performance may degrade when in-distribution data is noisy or poorly separated.
From a practical perspective, GROOD improves inference speed compared to KNN-based methods (see §˜A.4), but storing and processing gradients can be costly for very large datasets or high-resolution images. Prototype construction also matters: while relatively robust, performance can depend on the diversity of auxiliary OOD samples (§˜A.2).
Finally, several design choices proved important. Proximity-based filtering of synthetic prototypes (§˜5.2) enhanced discriminative power, and mixup toward the second-highest predicted class yielded stronger results than random interpolation, likely because it produces harder boundary samples.
In summary, GROOD offers a simple and effective framework for OOD detection, especially when in-distribution data is well-structured. Future work could extend it to noisier domains and explore ways to further reduce computational overhead.
Acknowledgments
This work was carried out within the DEEL project. It is supported by the DEEL Project CRDPJ 537462-18 funded by the Natural Sciences and Engineering Research Council of Canada (NSERC) and the Consortium for Research and Innovation in Aerospace in Québec (CRIAQ), together with its industrial partners Thales Canada inc, Bell Textron Canada Limited, CAE inc and Bombardier inc. [^1] https://deel.quebec. The DEEL project is also part of IRT Saint Exupéry and the ANITI AI cluster [^2] https://www.deel.ai/. The authors acknowledge the financial support from DEEL’s Industrial and Academic Members and the France 2030 program – Grant agreements n°ANR-10-AIRT-01 and n°ANR-23-IACL-0002. We are grateful for the digital alliance of Canada and NVIDIA for compute infrastructure. We would also like to thank Dr. Samer Nashed for his fruitful discussions and valuable feedback during the review process.
@@ A.1 GROOD Algorithm
For a comprehensive overview, the complete GROOD algorithm is outlined in algorithm˜1 and algorithm˜2.
@@
3:Compute class prototypes $\mathbf{P}_{\text{early}}$ and $\mathbf{P}_{\text{pen}}$ $\triangleright$ eq.˜7
4:Compute OOD prototype $\mathbf{p}^{\text{pen}}_{\text{ood}}$ using synthetic data generation§˜5.2 $\triangleright$ eq.˜8
@@
6:  compute $\nabla H(h)$ $\triangleright$ eq.˜5 using $\mathbf{p}^{\text{pen}}_{y=1,\cdots,C}$, $\mathbf{p}^{\text{pen}}_{\text{ood}}$
@@ A.2 Choice of OOD data for OOD prototype computation
To simulate OOD data representations, we utilize a manifold mixup technique on the early layer, similar to the targeted mixup approach described in §˜5. However, our method differs in the interpolation target. Instead of interpolating towards the predicted class prototype, we interpolate towards the second-highest predicted class $c_{2}$, which is the closest incorrect class on the decision boundary.
@@
Table˜6 illustrates our method’s robustness to different validation OOD datasets.
@@ A.3 Density Plots
To comprehensively evaluate the GROOD method’s ability to distinguish between in-distribution (ID) and out-of-distribution (OOD) data, we visualize the distribution of OOD scores across a range of datasets with varying characteristics. Figure˜4 presents these visualizations for ID, Near-OOD, and Far-OOD samples on CIFAR-10, CIFAR-100 (datasets with relatively small, natural images), ImageNet-200 (a subset of ImageNet), and ImageNet-1k (a large-scale, complex dataset). Analyzing performance across this spectrum demonstrates GROOD’s robustness to differences in image complexity and dataset size. In each subplot, we use density plots to represent the distribution of OOD scores, allowing for a clear visual comparison of separation between ID and OOD distributions.
@@ A.4 Inference Speed
GROOD introduces a novel Out-of-Distribution (OOD) detection method involving two forward passes for the mixup part which is inexpensive to compute, a backward pass over the OOD prototype which can be computed using its closed form expression as in eq.˜5 of the main paper and the ne arest neighbor search which is more computationally intensive. For the latter, Our approach employs the FAISS IndexIVF method for efficient distance computation, utilizing centroids and inverted lists instead of the complete dataset. This technique notably enhances inference speed compared to KNN, particularly in our CIFAR benchmarks. Specifically, on CIFAR-10 and CIFAR-100 datasets, GROOD recorded evaluation inference times over all OOD test sets of 130 seconds and 155 seconds, respectively. This is significantly faster than KNN, which took 434 seconds for CIFAR-10 and 641 seconds for CIFAR-100, demonstrating the efficiency of our approach.
@@ A.5 Ablate value of kth nearest neighbor
Figure˜5 illustrates the impact of using the distance to the $k$-th nearest neighbor, as proposed by Sun et al. (2022). The plot demonstrates that employing the distance to the nearest point in gradient space leads to optimal results.
@@ A.6 Gradient computation in closed form
This section details the derivation of the closed-form expression for the gradient, as presented in eq.˜5 of the main paper. The cross-entropy loss for a logit $L=(L_{j})_{j=1}^{C}+1$ and some (any) ID label $y$ is given by:
@@
which is equal to the softmax probability corresponding to the OOD class and does not depend on the specific ID class label $y$ anymore. Now for a feature vector $h$, the corresponding logit vector $L(h)$ is given by eq.˜2. Since $L_{C+1}(h)=\|h-\mathbf{p}^{\text{pen}}_{\text{ood}}\|_{2}$ is the only logit depending on $\mathbf{p}^{\text{pen}}_{\text{ood}}$, the gradient of the above loss with respect to $\mathbf{p}^{\text{pen}}_{\text{ood}}$ is given by the chain rule:
@@
A.7 Impact of Mixup-Trained Backbones on GROOD
We investigate how GROOD’s OOD prototype construction interacts with backbones trained using manifold mixup (Verma et al., 2019). Since these models are explicitly trained to generalize across interpolated samples, we hypothesized that synthetic mixup-based OOD prototypes might no longer serve as an effective deviation reference.
Method Near-OOD AUROC (%) Far-OOD AUROC (%)
GROOD (mean prototype) 81.05 80.26
ASH (Djurisic et al., 2022) 79.1 56.0
KNN (Sun et al., 2022) 78.0 81.85
GradNorm (Huang et al., 2021b) 50.0 50.0
Table 7: Performance of GROOD with mean prototype on a manifold mixup-trained ResNet-18 backbone (CIFAR-100).
To test this, we evaluated GROOD on a ResNet-18 trained with manifold mixup. Instead of using mixup-based OOD prototypes since it will no longer represent OOD data, we used a mean prototype computed from ID class prototypes. The results are shown below:
These results show that GROOD maintains competitive performance by adjusting its prototype strategy to the model’s training procedure. The mixup-based prototype remains optimal for standard-trained models, while alternative strategies like mean prototypes are preferable when the backbone is mixup-regularized.